OpenStudy (anonymous):

what is the exact area of a circle with a diameter of 37 inches

OpenStudy (lgbasallote):

\[Area = 3.14 \times (\frac{diameter}{2})^2\] \[\Rightarrow Area = 3.14 \times (\frac{37}{2})^2\]

OpenStudy (mertsj):

\[(\frac{37}{2})^2\pi\]

OpenStudy (mertsj):

The problem calls for the EXACT value so you must leave it in terms of pi